# Gaussian-Det: Learning Closed-Surface Gaussians for 3D Object Detection

### v1

@@ 1 Introduction
Figure 1: Top Left: Gaussians with color and coordinate. Top Right: Gaussians with color, coordinate and surface information, but in a much fewer amount. Bottom (a-b): Object composed of Gaussians exhibits a textured surface as opposed to the top left illustration. Bottom (c-d): Gaussian splatting inherently introduces numerous outliers around the object.
In this paper, we propose Gaussian-Det which exploits continuous surface representation for multi-view based 3D object detection. Gaussian-Det leverages the recently-proposed Gaussian Splatting [Kerbl et al., 2023] and formulates the input Gaussians into a mass of descriptors on partial surfaces. As shown in the top part of Figure 1, even with less than a third of the number, Gaussians utilizing surface information better demonstrate the texture information and showcase the smoothness and continuity. While it is feasible to directly employ point cloud based detectors, they can be distracted by numerous outliers inherently introduced by Gaussian Splatting process shown in Bottom (c-d) part of Figure 1. Moreover, the point-based detectors regard the input Gaussians as infinitesimal positional representations and discard the explicit depiction on continuous surface, which may lead to sub-optimal detection proposals. To address this, we design a Closure Inferring Module (CIM) for the comprehensive surface-based objectness deduction in Gaussian-Det. Specifically, taking the underdetermined nature of Gaussian Splatting into account, CIM firstly conducts the partial surface inference via estimating a variational term as a probabilistic feature residual to the original representation of partial surfaces. Under the assumption that the partial surfaces corresponding to an object can form a closed surface, CIM then coalesces them into a holistic representation which measures the closure of the overall proposal, serving as the prior on the quality and reliability of objectness prediction and the information basis of proposal refinement. To validate the effectiveness of the proposed Gaussian-Det, we have conducted experiments on both synthetic 3D-FRONT [Fu et al., 2021] and real-world ScanNet [Dai et al., 2017] datasets. The experimental results demonstrate that Gaussian-Det outperforms various existing approaches in terms of both average precision and recall.
@@ 2.1 Neural Scene Representation
In recent years, the methods of neural scene representation have presented astonishing performances on visual computing and geometry learning. They usually take a coordinate as input and output corresponding scene properties, such as signed distance [Park et al., 2019, Chabra et al., 2020], occupancy [Mescheder et al., 2019], color and density. Notably, in the scenario of multi-view geometry reconstruction, the Neural Radiance Field (NeRF) [Mildenhall et al., 2020] learns a plenoptic mapping from the querying coordinate and viewing direction into the point-wise radiation and volume density. The NeRF-based scene modeling has inspired numerous extended works spanning from low-level geometry reconstruction [Barron et al., 2021, Fridovich-Keil et al., 2022, Wang et al., 2021] to high-level visual understanding [Zhi et al., 2021, Jeong et al., 2022, Hu et al., 2023, Xu et al., 2023, Tian et al., 2024, Irshad et al., 2024]. Since NeRF requires a massive amount of queries during both training and inference phases, some improved representations are introduced as proxies such as hash grids [Müller et al., 2022], tri-planes [Chen et al., 2022] and points [Xu et al., 2022a]. Among these advancements, the 3D Gaussian Splatting (3D-GS) [Kerbl et al., 2023] stands out for its balance between quality and efficiency. 3D-GS employs a set of anisotropic Gaussians initialized by Structure from Motion (SfM), whose amount and properties are iteratively updated during training. Apart from the positional information as in point cloud based or NeRF-based representation, a 3D Gaussian naturally depicts a continuous surface through its mean, covariance parameters. Based on 3D-GS, several works have explored on 3D content generation [Tang et al., 2024], surface reconstruction [Guédon and Lepetit, 2024], semantic segmentation [Shi et al., 2024] and editing [Chen et al., 2024].
2.2 Indoor 3D Object Detection
@@
Indoor 3D object detection has drawn extensive attention owing to its broad applications in robotic vision, augmented reality, etc. The techniques can be categorized according to the input modalities, such as point cloud, gridded voxels and 2D image(s). For example, Song and Xiao [2016] uses a volumetric CNN to create 3D Region Proposal Network (RPN) on a voxelized 3D scene. VoteNet [Qi et al., 2019] directly operates on point cloud by using a PointNet++ [Qi et al., 2017b] backbone and center voting strategy. The follow-up point-based approaches include the VoteNet-style  [Cheng et al., 2021, Liu et al., 2021] and the sparse convolution based methods [Rukhovich et al., 2022a, Wang et al., 2022]. However, the reliance on expensive lidars or depth sensors limits their applicability. In contrast, the RGB camera systems offer a cost-effective alternative. To overcome the inherent ill-posedness in image-based 3D object detection, ImVoxelNet [Rukhovich et al., 2022b] lifts the 2D features into a 3D volume representation followed by a FCOS [Tian et al., 2019]-style detection head. ImGeoNet [Tu et al., 2023] introduces an image-induced geometry-aware voxel representation for better preservation of geometry. However, the projective geometry extracted by Rukhovich et al. [2022b], Tu et al. [2023] still originates from 2D features and may thus result in ambiguous detections. On the other hand, several advances [Hu et al., 2023, Xu et al., 2023, Irshad et al., 2024] seek to leverage the multi-view consensus and consistency of NeRF representation [Mildenhall et al., 2020] that embeds the geometry into the voxel grids. NeRF-RPN consumes the volumetric features sampled from NeRF and then directly regresses the bounding boxes via a generic RPN. NeRF-Det learns a two-branch framework that simultaneously conducts geometry reconstruction and bounding box regression. These two approaches rely on uniform and discrete sampling which may fail to capture the true objectness in 3D scenes. In contrast, our proposed Gaussian-Det strives for representing the targets in a continuous manner by leveraging the descriptors on partial surfaces and coalescing them into a holistic measurement which serves as a comprehensive prior on the reliability of the deducted objectness.
@@ 3 Approach
Figure 2: Framework of our Gaussian-Det. Gaussian-Det takes as input the Gaussians reconstructed from scanned multi-view images of an indoor scene. We formulate surface-based Gaussian representation which is firstly used to predict the initial object proposal. Then the closure inferring module, which contains partial surface feature inference and holistic surface closure coalescence, learns a reliability measurement of the proposal and estimates the refined detection result, whose absolute closure score is low similar to the ground truth.
@@ 3.2 Surface-based Gaussian Representation and Proposal Initialization
The reformulated Gaussian representation $\mathbb{G}^{surf}$ initializes the depiction on multiple partial surfaces. Given the huge amount of Gaussians in $\mathbb{G}^{surf}$, we use an off-the-shelf backbone $\mathcal{B}(\cdot)$ for trimming:
@@
where $\mathbb{G}^{cand}=\{\mathbf{g}_{1},\mathbf{g}_{2},\dots,\mathbf{g}_{M}\}\in\mathbb{R}^{M\times 7}$ is a subset of $\mathbb{G}^{surf}$ and $\mathcal{F}^{cand}=\{\mathbf{f_{1}},\mathbf{f_{2}},\dots,\mathbf{f_{M}}\}\in\mathbb{R}^{M\times C}$ denotes the corresponding feature in the latent space. From $\mathbb{G}^{surf}$, we then accordingly initialize the partial-aware object proposals $\mathbf{P}^{part}=\{p_{1},p_{2},\dots,p_{K}\}$ by employing a VoteNet [Qi et al., 2019]-style head [Cheng et al., 2021] which groups $\mathbb{G}^{cand}$ in Euclidean space and predicts the bounding boxes with a MLP-based architecture [Qi et al., 2017a]. Meanwhile, the partial-aware features of each proposal is extracted:
$\mathbf{f}^{part}_{k}=\mathcal{M}(\mathcal{F}^{cand},p_{k})$ (5)
where $\mathcal{M}$ is a querying-and-pooling function [Cheng et al., 2021] that operates around the location of each proposal.
@@ 3.3 Closure Inference Module for Proposal Refinement
As illustrated in Figure 2, the Closure Inference Module (CIM) consists of two stages: partial surface feature inference and holistic surface closure coalescence.
Partial Surface Feature Inference: CIM firstly conducts the partial surface modeling via probabilistic feature construction. In 3D Gaussian Splatting, inferring local surface features for spatial points is underdetermined due to the supervision using 2D signals, leading to the unstructured outliers [Guédon and Lepetit, 2024] with unreliable local surface descriptors. To this end, we introduce a variational term $\mathcal{F}^{V}$ into the candidate features $\mathcal{F}^{cand}$, establishing a probabilistic feature residual:
@@
where $\alpha$ is a weight hyper-parameter (in practice we set $\alpha=0.1$). The randomness is modeled by assuming $\mathcal{F}^{V}$ can be sampled from $\mathbf{p}_{\theta}(\mathcal{F}^{V})=N(\mathcal{F}^{V};\mathbf{0},\mathbf{I})$. To fit the posterior $\log q_{\phi}(\mathcal{F}^{V}|\mathcal{F}^{cand})$ and likelihood $p_{\theta}(\mathcal{F}^{cand}|\mathcal{F}^{V})$ parameterized by $\phi$ and $\theta$ respectively, the following evidence lower bound (ELBO) is optimized for the residual estimation:
@@
Figure 3: Example of utilizing closure as a prior for determining the quality of objectness prediction.
Holistic Surface Closure Coalescence: The partial surfaces are then coalesced into a holistic representation by CIM, which provides informative clues on surface closure for objectness deduction. As illustrated in Figure 3, we assume that the partial surfaces corresponding to an object can form a closed surface, which serves as a prior for determining the quality of objectness predictions. The prior can be mathematically expressed by Theorem 1.
@@
Proof.
It is a special case of Gauss’ Flux Theorem by setting $\mathbf{T}$ as any constant vector field:
$\oiint\limits_{S}\mathbf{T}\cdot\mathbf{n}\mathrm{d}S=\iiint\limits_{V}\nabla\cdot\mathbf{T}\mathrm{d}V\stackrel{{\scriptstyle\nabla\cdot\mathbf{T}=0}}{{=}}0\qed$ (9)
Numerically we implement the integrated flux in Eqn. (9) for the $k$-th proposal as quadrature:
@@
With the estimated flux as the measurement of surface closure, we formulate the geometrically coalesced representation as:
$\mathbf{f}^{hol}_{k}=S_{k}\cdot\mathcal{M}(\mathcal{F}^{cand},p_{k}),\;\text{where}\;S_{k}=\exp(-\gamma_{k}|\hat{\mathbf{\Phi}}_{k}|)$ (11)
where $\mathcal{F}^{cand}$ is the inferred features with the probabilistic residual in Eqn.(6). $S_{k}$ serves as the closure weight parameter. $\gamma_{k}$ is a flexible coefficient mapped from $\mathcal{F}^{V}$ through a MLP and sigmoid activation function. In this way, it controls the preservation (when $\hat{\mathbf{\Phi}}_{k}$ is close to zero) or suppression (when the absolute value of $\hat{\mathbf{\Phi}}_{k}$ is very large) of $\mathbf{f}^{hol}_{k}$.
We finally formulate the refined proposal feature which comprehensively combines the information on partial-aware surfaces and prior on holistic closure-based measurement on objectness as:
@@
where [||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] denotes the concatenation operation along the feature dimension. $\mathbf{f}^{final}_{k}$ is used for proposal refinement by estimating a parameter offset to the original proposals through a stacked MLP:
@@ 4.1 Experimental Settings and Implementation Details
We experimented Gaussian-Det on the officially-released benchmark by Hu et al. [2023] which contain both the synthetic 3D-FRONT [Fu et al., 2021] and the real-world ScanNet [Dai et al., 2017] datasets. We faithfully followed Hu et al. [2023] to take the posed images as input. The camera poses in the datasets released by Hu et al. [2023] were estimated via Structure from Motion (SfM) which used the groundtruth point cloud as initialization.
The synthetic 3D-FRONT [Fu et al., 2021] dataset is featured with its large-scale room layouts and textured furniture models. The complex scenes are manually downsized into 159 usable rooms after data cleansing by Hu et al. [2023]. The indoor scenes are densely annotated with oriented bounding boxes. We follow Hu et al. [2023] to remove the annotations of construction objects such as ceilings and floors, and merge the components that belong to the same object instance into a whole.
The real-world ScanNet [Dai et al., 2017] dataset consists of complex and challenging scenes captured by Structure Sensor in an ego-centric manner. We followed Hu et al. [2023] to uniformly divide the frames into 100 bins and choose the sharpest one in each bin measured by Laplacian variance. Experiments were done on the 90 scenes that are randomly selected by Hu et al. [2023].
For the input Gaussians $\mathbb{G}$, we used the official implementation of 3D-GS [Kerbl et al., 2023] equipped with the SuGaR regularization [Guédon and Lepetit, 2024], which takes 30,000 training iterations. We excluded the Gaussians outside the room cuboid defined by Hu et al. [2023] and filtered out Gaussians with opacity value lower than 0.3. Note that such a combined implementation [Kerbl et al., 2023, Guédon and Lepetit, 2024] aims at decent 2D rendering results. The issue of outlier Gaussians around the 3D objects still severely exists as shown in Figure 1. We employed PointNet++ [Qi et al., 2017b] as the backbone $\mathcal{B}$. We used a single NVIDIA RTX A6000 48GB GPU for both training and evaluation. The total training epochs of detection is 50 for both 3D-FRONT and ScanNet. The evaluation metrics are mainly average recall (AR) and average precision (AP), thresholded by IoU value at both 0.25 and 0.5. Please refer to the Appendix pages for the details on model hyper-parameters.
Figure 4: Statistics on surface closure of two datasets [Fu et al., 2021, Dai et al., 2017]. Most objects in both datasets contain a higher degree of surface closure. The statistics are collected from our reconstructed 3D Gaussians.

### v2

@@ Abstract
Figure 1: We leverage Gaussian Splatting as surface representation for multi-view based 3D object detection. Top Left: Gaussians with color and coordinate. Top Right: Gaussians with color, coordinate and surface information, but in a much fewer amount. Bottom (a-b): Object composed of Gaussians exhibits a textured surface as opposed to the top left illustration. Bottom (c-d): However, Gaussian splatting inherently introduces numerous outliers around the object.
@@ 1 Introduction
Figure 2: Illustration of the surface closure prior in Gaussian-Det. For 3D Gaussians within a bounding box, those forming relatively closed surfaces (lower $|\Phi|$) indicate an accurate detection, and vice versa. Note that 3D Gaussians outside the bounding box are not shown for clarity.
@@
In this paper, we propose Gaussian-Det which exploits continuous surface representation for multi-view based 3D object detection. Gaussian-Det leverages the recent Gaussian Splatting [Kerbl et al., 2023] and formulates the input Gaussians into a mass of descriptors on partial surfaces. As shown by top of Figure 1, even with less than a third of the number, Gaussians utilizing surface information better demonstrate the texture information, smoothness and continuity. While it is feasible to directly employ point cloud based detectors, they can be distracted by numerous outliers inherently introduced by Gaussian Splatting process shown in Bottom (c-d) part of Figure 1. Moreover, the point-based detectors regard the input Gaussians as infinitesimal positional representations and discard the explicit depiction on continuous surface, which may lead to sub-optimal detection proposals. To address this, we design a Closure Inferring Module (CIM) for the comprehensive surface-based objectness deduction in Gaussian-Det. As displayed in Figure 2, CIM is built upon the assumptive prior that, the partial surfaces in an accurately detected bounding box can form a relatively closed surface measured by a lower flux value. Specifically, taking the underdetermined nature of Gaussian Splatting into account, CIM firstly conducts the partial surface inference via estimating a variational term as a probabilistic feature residual to the original representation of partial surfaces. CIM then coalesces them into a holistic representation and uses the flux value measures the closure of the overall proposal, serving as the prior on the quality and reliability of objectness prediction and the information basis of proposal refinement. To validate the effectiveness of the proposed Gaussian-Det, we have conducted experiments on both synthetic 3D-FRONT [Fu et al., 2021] and real-world ScanNet [Dai et al., 2017] datasets. The experimental results demonstrate that Gaussian-Det outperforms various existing approaches in terms of both average precision and recall.
@@ 2 Related Work
Neural Scene Representation: In recent years, neural scene representation have presented astonishing performances on visual computing and geometry learning. They usually take a coordinate as input and output corresponding scene properties, such as signed distance [Park et al., 2019, Chabra et al., 2020], occupancy [Mescheder et al., 2019], color and density. Notably, in the scenario of multi-view geometry reconstruction, the Neural Radiance Field (NeRF) [Mildenhall et al., 2020] learns a plenoptic mapping from the querying coordinate and viewing direction into the point-wise radiation and volume density. The NeRF-based scene modeling has inspired numerous extended works spanning from low-level geometry reconstruction [Barron et al., 2021, Fridovich-Keil et al., 2022, Wang et al., 2021] to high-level visual understanding [Zhi et al., 2021, Jeong et al., 2022, Hu et al., 2023, Xu et al., 2023, Liu et al., 2023, Tian et al., 2024, Irshad et al., 2024]. Since NeRF requires a massive amount of queries during both training and inference phases, some improved representations are introduced as proxies such as hash grids [Müller et al., 2022], tri-planes [Chen et al., 2022] and points [Xu et al., 2022a]. Among these advancements, the 3D Gaussian Splatting (3D-GS) [Kerbl et al., 2023] stands out for its balance between quality and efficiency. 3D-GS employs a set of anisotropic Gaussians initialized by Structure from Motion (SfM), whose amount and properties are iteratively updated during training. Apart from the positional information as in point cloud based or NeRF-based representation, a 3D Gaussian naturally depicts a continuous surface through its mean, covariance parameters. Based on 3D-GS, several works have explored on novel view synthesis Charatan et al. [2024], Chen et al. [2024b], Zhang et al. [2024], 3D content generation [Tang et al., 2024], surface reconstruction [Guédon and Lepetit, 2024], semantic segmentation [Shi et al., 2024] and editing [Chen et al., 2024a].
@@
Indoor 3D Object Detection: Indoor 3D object detection has drawn extensive attention owing to its broad applications in robotic vision, augmented reality, etc. The techniques can be categorized according to the input modalities, such as point cloud, gridded voxels and 2D image(s). For example, [Song and Xiao, 2016] uses a volumetric CNN to create 3D Region Proposal Network (RPN) on a voxelized 3D scene. VoteNet [Qi et al., 2019] directly operates on point cloud by using a PointNet++ [Qi et al., 2017b] backbone and center voting strategy. The follow-up point-based approaches include the VoteNet-style  [Cheng et al., 2021, Liu et al., 2021, Zheng et al., 2024] and the sparse convolution based methods [Rukhovich et al., 2022a, Wang et al., 2022]. However, the reliance on expensive lidars or depth sensors limits their applicability. In contrast, the RGB camera systems offer a cost-effective alternative. To overcome the inherent ill-posedness in image-based 3D object detection, ImVoxelNet [Rukhovich et al., 2022b] lifts the 2D features into a 3D volume representation followed by a FCOS [Tian et al., 2019]-style detection head. ImGeoNet [Tu et al., 2023] introduces an image-induced geometry-aware voxel representation for better preservation of geometry. However, the projective geometry extracted by [Rukhovich et al., 2022b, Tu et al., 2023] still originates from 2D features and may thus result in ambiguous detections. On the other hand, several advances [Hu et al., 2023, Xu et al., 2023, Irshad et al., 2024] seek to leverage the multi-view consensus and consistency of NeRF representation [Mildenhall et al., 2020] that embeds the geometry into the voxel grids. NeRF-RPN consumes the volumetric features sampled from NeRF and then directly regresses the bounding boxes via a generic RPN. NeRF-Det learns a two-branch framework that simultaneously conducts geometry reconstruction and bounding box regression. These two approaches rely on uniform and discrete sampling which may fail to capture the true objectness in 3D scenes. In contrast, our proposed Gaussian-Det strives for representing the targets in a continuous manner by leveraging the descriptors on partial surfaces and coalescing them into a holistic measurement which serves as a comprehensive prior on the reliability of the deducted objectness.
@@ 3 Approach
[…]
@@ 3.2 Surface-based Gaussian Representation and Proposal Initialization
The reformulated Gaussian representation $\mathbb{G}^{surf}$ initializes the depiction on multiple partial surfaces. Given the huge amount of Gaussians in $\mathbb{G}^{surf}$, we use an off-the-shelf backbone $\mathcal{B}(\cdot)$ for simultaneously trimming them into a candidate subset and extracting their corresponding features:
@@
where $\mathbb{G}^{cand}=\{\mathbf{g}_{1},\mathbf{g}_{2},\dots,\mathbf{g}_{M}\}\in\mathbb{R}^{M\times 7}$ is a subset of $\mathbb{G}^{surf}$ and $\mathcal{F}^{cand}=\{\mathbf{f_{1}},\mathbf{f_{2}},\dots,\mathbf{f_{M}}\}\in\mathbb{R}^{M\times C}$ denotes the corresponding feature in the latent space. From $\mathbb{G}^{cand}$, we then accordingly initialize the partial-aware object proposals $\mathbf{P}^{part}=\{p_{1},p_{2},\dots,p_{K}\}$ by employing a VoteNet [Qi et al., 2019]-style head [Cheng et al., 2021] which groups $\mathbb{G}^{cand}$ in Euclidean space and predicts the bounding boxes with a MLP-based architecture [Qi et al., 2017a]. Meanwhile, the partial-aware features of each proposal is extracted:
$\mathbf{f}^{part}_{k}=\mathcal{M}^{part}(\mathcal{F}^{cand},p_{k})$ (5)
where $\mathcal{M}^{part}$ is a querying-and-pooling function [Cheng et al., 2021] that operates around the location of each proposal.
@@ 3.3 Closure Inference Module for Proposal Refinement
As illustrated in the bottom part of Figure 3, the Closure Inference Module (CIM) consists of two stages: Given the underdetermined nature of 3D Gaussian Splatting, CIM firstly enhances the representation of partial surfaces via variationally estimating the probabilistic feature residuals. Subsequently, under the assumption that the higher degree of surface closure indicates a more accurate detection (and vice versa), CIM seeks to a holistic surface closure coalescence, serving as an informative prior on the quality of the predicted objectness.
Partial Surface Inference: In Gaussian Splatting, inferring local surface features for spatial points is underdetermined due to the supervision from 2D signals, leading to unstructured outliers [Guédon and Lepetit, 2024] with unreliable local surface descriptors. To this end, we model such randomness in CIM by incorporating a variational term $\mathcal{F}^{V}$ into the candidate features $\mathcal{F}^{cand}$, establishing a probabilistic feature residual:
@@
where $\alpha$ is a weight hyper-parameter (in practice we set $\alpha=0.1$). The randomness is modeled by assuming $\mathcal{F}^{V}$ can be sampled from $\mathbf{p}_{\theta}(\mathcal{F}^{V})=N(\mathcal{F}^{V};\mathbf{0},\mathbf{I})$. To fit the posterior $\log q_{\phi}(\mathcal{F}^{V}|\mathcal{F}^{cand})$ and likelihood $p_{\theta}(\mathcal{F}^{cand}|\mathcal{F}^{V})$ parameterized by $\phi$ and $\theta$ respectively, the following evidence lower bound (ELBO) is optimized for the residual estimation[Kingma and Welling, 2014]:
@@
Holistic Surface Closure Coalescence: The partial surfaces are then coalesced into a holistic representation by CIM. Specifically, as illustrated in Figure 2, CIM leverages a key geometric property that the partial surfaces corresponding to an object can approximate a closed surface. As formulated by Theorem 1 (see Appendix for proof), this property can be mathematically expressed that the net flux $\phi$ of a constant vector field through a closed surface is zero.
@@
In CIM, we utilize this theorem by computing the flux of a pre-set constant vector field over the partial surfaces enclosed by an object proposal. The closer the total flux is to zero, the better the surfaces form a closed object, suggesting a higher-quality detection. To practically implement this concept for the $k$-th object proposal, CIM computes the estimated flux $\hat{\bm{\Phi}}_{k}$ numerically using quadrature:
@@
With the estimated flux as the quantitative measurement of surface closure, CIM then formulate the geometrically coalesced representation as:
$\mathbf{f}^{hol}_{k}=S_{k}\cdot\mathcal{M}^{hol}(\mathcal{F}^{cand},p_{k}),\;\text{where}\;S_{k}=\exp(-\gamma_{k}|\hat{\mathbf{\Phi}}_{k}|)$ (10)
where $\mathcal{F}^{cand}$ is the enhanced feature with the probabilistic residual in Eqn.(6). $\gamma_{k}$ is a flexibility coefficient mapped from $\mathcal{F}^{V}$ through a MLP and sigmoid activation function. $\mathcal{M}^{hol}$ is another querying-and-pooling function sharing the same architecture with $\mathcal{M}^{part}$. $S_{k}$ serves as the closure weight parameter which controls the support ($|\hat{\mathbf{\Phi}}_{k}|\downarrow,S_{k}\uparrow$) or suppression ($|\hat{\mathbf{\Phi}}_{k}|\uparrow,S_{k}\downarrow$) of $\mathbf{f}^{hol}_{k}$.
We finally formulate the refined proposal feature which comprehensively combines the information on partial-aware surfaces and holistic closure-based measurement on objectness as:
@@
where [||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] denotes the concatenation along the feature dimension. $\mathbf{f}^{final}_{k}$ is used for proposal refinement by estimating a parameter offset to the original proposals through a stacked MLP:
@@ 4.1 Experimental Settings and Implementation Details
We experimented Gaussian-Det on the officially-released benchmark by [Hu et al., 2023] which contain both the synthetic 3D-FRONT [Fu et al., 2021] and the real-world ScanNet [Dai et al., 2017] datasets. We faithfully followed [Hu et al., 2023] to take the posed images as input. The camera poses in the datasets released by [Hu et al., 2023] were estimated via Structure from Motion (SfM) which used the groundtruth point cloud as initialization.
The synthetic 3D-FRONT [Fu et al., 2021] dataset is featured with its large-scale room layouts and textured furniture models. The dataset contains 159 usable rooms that are manually selected, cleaned and rendered by [Hu et al., 2023]. The indoor scenes are densely annotated with oriented bounding boxes. We follow [Hu et al., 2023] to remove the annotations of construction objects such as ceilings and floors, and merge the components that belong to the same object instance into a whole. Please refer to the Appendix pages for illustrative examples of this dataset.
The real-world ScanNet [Dai et al., 2017] dataset consists of complex and challenging scenes captured by Structure Sensor in an ego-centric manner. We followed [Hu et al., 2023] to uniformly divide the frames into 100 bins and choose the sharpest one in each bin measured by Laplacian variance. Experiments were done on the 90 scenes that are randomly selected by [Hu et al., 2023].
For the input Gaussians $\mathbb{G}$, we used the official implementation of 3D-GS [Kerbl et al., 2023] equipped with the SuGaR regularization [Guédon and Lepetit, 2024], which takes 30,000 training iterations. We excluded the Gaussians outside the room cuboid defined by [Hu et al., 2023] and filtered out Gaussians with opacity value lower than 0.3. Note that such a combined implementation [Kerbl et al., 2023, Guédon and Lepetit, 2024] aims at decent 2D rendering results. The issue of outlier Gaussians around the 3D objects still severely exists as shown in Figure 1. We employed PointNet++ [Qi et al., 2017b] as the backbone $\mathcal{B}$. We used a single NVIDIA RTX A6000 48GB GPU for both training and evaluation. The total training epochs of detection is 50 for both 3D-FRONT and ScanNet. The evaluation metrics are mainly average recall (AR) and average precision (AP), thresholded by IoU value at both 0.25 and 0.5. Please refer to the Appendix pages for the details on model hyper-parameters.
From the statistical results (see Appendix) of our reconstructed 3D Gaussians on two datasets, we observe that the majority of flux values are clustered near zero. This indicates that most objects contain a higher degree of surface closure and our assumption is applicable to the employed datasets.